Hawaii-2017-05 experiment with new MOBYrefresh - BS03cfg002 & 003 for MOBY263 deployment

Iris check and system responses. Using Iris's and neutral density filters to reduce Es/Ed 100x and Lu by 10x

Trk # 1 = Es
Trk # 3 = EdTop
Trk # 5 = EMid
Trk # 7 = EdBot
Trk # 9 = LuTop
Trk #11 = LuMid
Trk #13 = LuBot

BS03cfg002 = pre-L264, 29-Jul-2017 = shutter block & iris installed (no shutter controller)
BS03cfg003 = pre-L264, 31-Jul-2017 = irises glued down, shutter controller installed
Problems seen (and solved) and good things:
  1. Mike found during cals the shutter control on BS03 return 5V = open for track 10 when Track 9 is open and the reverse. The correct shutter opens it just returns the track is open on the wrong track.
  2. No UV ghosting found

Log Sheets

Photos

 

All the raw data files and their images , Table of KEYWORDS variable
Page Number
Link
Description
Date
1.01
Day 01 - 29 Jul Raw Data , Table
DAY01 - Iris and neutral density filter track reduction check - cfg002 -> shutter block & iris installed (no shutter controller)
29 Jul 2017
1.02
Day 02 - 31 Jul Raw Data , Table
DAY02 - Lu system responses - cfg003 -> irises glued down, shutter controller installed
31 Jul 2017
1.03
Day 03 - 1 Augl Raw Data , Table
DAY03 - Es system response
31 Jul 2017
 
Day01 - Iris and neutral density filter track reduction check
2.01 Iris Check Iris check for Es, Ed and Lu (Es/Ed 100x and Lu 10x) 31 Jul 2017
Day02 - Lu system responses
3.01 Lu system response Lu system responses 31 Jul 2017
Day03 - Es system response
4.01 findtracks Find tracks from day 1-3 1 Aug 2017
4.02 Es rsp Es system response 1 Aug 2017
4.03 Ghosting check Checking to see if light from track 1 is showing up on track 14 in the UV 19 Jan 2018
       
       

 

pre-experiment EMAIL from Mike

HI Mark,
R.E. adding iris to the BS03 to be deployed in MOBY263,
which will have sensors = 3x Lu + 3x Ed + Es:

It looks like we put irises on all 4x ports which were used from BS01 for MOBY262 = Es + 3x Lu.
I have notes noting that we wanted "15x, 22x, 15x less sensitive",
which would have been set during HI-2017-01, Day01,
which I think refer to the LuTop, Mid, Bot sensors (?).

And, I see some email from Dec-2016 ... Feb-2017 r.e. predicted exposure time,
from which I think the attenuation targets must have come (?).

The details are very fuzzy for me right now...
Maybe we can get some help from Steph et. al. tomorrow during the telecon!
MF

DAY 1 EMAIL from Mike

Hi Steph,
Mark assembled much of the BS03 today, and I just took some data that I was hoping you could look at please. I tried to adjust the irises to get a 10x reduction in counts for the Lu tracks and a 100x reduction for the Es & Ed tracks.  These reductions were versus port #14 as a reference, where #14 had no iris and no filter(s).

The BS03cfg02 setup is:
Trk #1=Es, #3=Et, #5=Em, #7=Eb, #9=Lt, #11=Lm, #13=Lb,
all of these tracks have an iris mounted next to the shutter,
and, the Es track also has 2x Nuetral Density filters = 1.0 OD + 0.5 OD,
and, the Ed tracks also have 1x ND filter = 1.0 OD (OD = Optical Density).
The filters are Thorlabs Reflective 012.7 mm NDUV Filter(s)
OD=1.0 == NDUV510B,  OD=0.5 == NDUV505B
(I took a photo of the filter boxes for documentation).

There are .fits data at your NAS/FTP, /ftp1/Mike/L264/bscal and some photos at .../photos, and 3x log sheets at .../doc. All files except the last #19 had trk #14 open, as a reference. File#19 had trk #14 closed, and the exposure time was increased to get good signal levels at the Lu tracks.

Pg.1 of the logsheet doc shows my confused mind at the start: I was not accounting for the background count level whilst trying to get the proper iris-reduced signals.  Pg.2 is the data that I think should be OK, starting at file #10.fits.  The files #5,6,7 from pg.1 should be interesting to look at though, because I completely closed the irises for the Ed tracks - it will be interesting to see how much light gets through a closed iris...

Thank You, MF

DAY 2 EMAIL from Mike

Hi Steph,
Mark got the shutter controller into the BS today, and he expoxied down the adjusted irises, while Sean & I got some of the BS fibers connected at
the bottom fiber block of the BS.  There are some data for you to look at, at /ftp1/Mike/L264/bscal = OL425(L9) scans via BS03cfg03 LuBot, Mid, Top.
These were collected via Andor SOLIS software via USB, and are .fits files, where s*=sig=open shutter (one shutter open at a time), d*=drk=shutter(s) closed,
and b*=bac=internal Andor camera scans (i.e. not firing the Andor shutter). There is 1x data sheet, which includes the foot-Lambert readout for the OL425.
I have LOTS of photos, but they are not organized yet...

I'm hoping tomorrow to get FEL scans via Es, and preferably a set of exposure time cals via the NIST OL455 sphere & LuBot.  The Ed's will not get pre-cal FEL scans,
on account of the arms had to be bolted/taped/painted and the GS5000 50 cm reference cannot get at Ed with the arms like this.

Thanks eh, MF

DAY 3 EMAIL from Mike

HI sTePh,
There are some more data for you at /ftp1/Mike/L264/bscal !
Only 5x sig,drk pairs from Es via FEL F601(C3), which is the same lamp+cal used on recent odd MOBY MOS E* sensors.  The exposure time was 4 min, so it took lots of time to get just 10 scans.  I forgot to cover the BS fiber bundle area of the BS during the first 3x scans, but there was a black cloth over said bundle during scans #4 & 5, so it will be interesting to see if 1:3 differ from 4:5. We ran out of time for an exposure time cal - so sorry.  Prior to the Es cal I would not have thought to run an exposure cal out to/past 240 sec...
THankS agaIn, MF

P.S. I found this for the FEL <attached: F601c3.txt>